# De-mark: Watermark Removal in Large Language Models

### v1

@@ 5.2 Effectiveness of De-mark
TPR@FPR= median p-value
0.1% 0.01% 0.001%
@@ 6 Conclusion
Appendix A Additional algorithms
Algorithm 4 Calculate relative probability ratio
2:Output: relative ratio matrix $R\in\mathbb{R}^{m\times m}$
3:Initialize $R=0$
4:for $i=1,\dots,m$ do
5:     for $j=i+1,\dots,m$ do
6:         $P(T_{i}|\bm{x},T_{i},T_{j}),P(T_{j}|\bm{x},T_{i},T_{j})=F(\bm{x},T_{i},T_{j})$
7:         $R_{i,j}=log(\frac{P(T_{i}|\bm{x},T_{i},T_{j})}{P(T_{j}|\bm{x},T_{i},T_{j})})$
8:         $R_{j,i}=log(\frac{P(T_{j}|\bm{x},T_{i},T_{j})}{P(T_{i}|\bm{x},T_{i},T_{j})})$
9:Return: $R$
Algorithm 5 Calculate token score
2:Output: token scores $\bm{s}\in\mathbb{R}^{m}$
3:Initialize $\bm{s}=0$
4:Generate $R$ using $\bm{x}$, $T$ with Alg.4
5:for $i=1,\dots,m$ do
6:     for $j=1,\dots,m$ do
7:         if $\alpha_{1}<R_{i,j}<\alpha_{2}$ then
8:              $s_{i}+=1$
9:         else if $-\alpha_{2}<R_{i,j}<-\alpha_{1}$ then
10:              $s_{i}-=1$
11:Return: $\bm{s}$
Appendix B Discussion
@@
Appendix C Experimental setting details
C.1 Implementation of probability estimation function and prompts
@@
Randomly start your answer with: "[x] [T_i]" or "[x] [T_j]"
@@
response on a scale of 1 to 100 by strictly following this format: "[[rating]]", for example:
"Rating: [[51]]".
@@
C.2 Main experiment settings
@@
For De-mark  hyperparameters, we use $\alpha_{1}=0.2$, $\alpha_{2}=10$, $\beta=0.8$, $\gamma=0.1$. The target token size $m$ in Alg.1 and Alg.2 is set to 50. The repeat time $c$ in Alg.2 is set to 5. We also found that De-mark is not sensitive to these hyperparameters.
@@
Appendix D Ablation study
@@
TPR@FPR= median p-value GPT Score
0.1% 0.01% 0.001%
@@
Appendix E Additional experimental results
@@
TPR@FPR= median p-value GPT score
0.1% 0.01% 0.001%
@@
Appendix F Proof of Theorem 4.1
@@
Assume w.l.o.g. $T_{i}$ in the green list and $T_{j}$ in the red list. Based on Equation 1, the watermarked distribution $P_{W}(T_{i}|\bm{x},T_{i},T_{j})$ and $P_{W}(T_{j}|\bm{x},T_{i},T_{j})$ should be $P_{W}(T_{i}|\bm{x},T_{i},T_{j})=\frac{(0.5+\epsilon)e^{\delta}}{(0.5+\epsilon)e^{\delta}+(0.5-\epsilon)}$ and $P_{W}(T_{j}|\bm{x},T_{i},T_{j})=\frac{(0.5-\epsilon)}{(0.5+\epsilon)e^{\delta}+(0.5-\epsilon)}$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ is the noise. In this case,
@@
Appendix G Proof of Theorem 4.2
@@
Appendix H Broader Impact Statement
Machine learning models are transforming various sectors by improving efficiencies and tackling complex problems (Yang et al., 2020, 2019; Wen et al., 2023; Chakraborty et al., 2022; Cai et al., 2022; Chen et al., 2024a; Xu and Li, 2017; Liu et al., 2024; Wen et al., 2024). Despite these beneficial impacts, there are valid concerns regarding the integrity and security of machine learning implementations (Hong et al., 2024; Hu et al., 2023b; Wang et al., 2023c, b; Wu et al., 2022, 2023a, 2023c). In this context, watermarking has become a critical tool. It not only verifies the authenticity and ownership of digital media but also aids in identifying AI-generated content.

### v2

@@ 5.2 Effectiveness of De-mark
TPR@FPR= median $\uparrow$
0.1$\downarrow$ 0.01$\downarrow$ 0.001$\downarrow$
@@ 6 Conclusion
Impact Statement
Acknowledgments
This work was partially supported by NSF IIS 2347592, 2348169, DBI 2405416, CCF 2348306, CNS 2347617.
Appendix A Discussion
@@
Appendix B Experimental setting details
B.1 Implementation of probability estimation function and prompts
@@
Randomly start your answer with: ”[x] [T_i]” or ”[x] [T_j]”
@@
response on a scale of 1 to 100 by strictly following this format: ”[[rating]]”, for example:
”Rating: [[51]]”.
@@
B.2 Main experiment settings
@@
For De-mark  hyperparameters, we use $\alpha_{1}=0.2$, $\alpha_{2}=10$, $\beta=0.8$, $\gamma=0.1$. The target token size $m$ in Alg.3 and Alg.4 is set to 50. The repeat time $c$ in Alg.4 is set to 5. We also found that De-mark is not sensitive to these hyperparameters.
@@
Following the previous work (Jovanović et al., 2024), we use three datasets for our experiments, Dolly creative writing (Conover et al., 2023), MMW Book Report (Piet et al., 2023), and MMW Story (Piet et al., 2023), each contains about 100 prompts for open-end writing. We also include additional experiments on WaterBench (Tu et al., 2023), selecting two tasks (Entity Probing and Concept Probing) with short outputs, and one task (Long-form QA) with long outputs. The results are presented in table 7. The findings suggest that our proposed De-mark method performs well across these tasks. Furthermore, when the output is short, watermarking has a more significant impact on output quality, and our method effectively recovers the original distribution, thereby improving performance.
Table 7: Additional experimental results on watermark removal task.
TPR@FPR=
Entity Probing (short answer) 10% ↓ 1% ↓ 0.1% ↓ Median p-value ↑ F1 ↑
Llama3.1 8B +Watermark 51.5% 18.5% 9.0% 8.07e-2 12.49
Llama3.1 8B +DE-MARK 14.5% 1.5% 0.0% 3.69e-1 12.03
Mistral 7B +Watermark 50.0% 13.5% 3.5% 9.94e-2 7.46
Mistral 7B +DE-MARK 16.0% 2.5% 0.5% 3.60e-1 7.64
Llama3.2 3B +Watermark 40.0% 15.5% 6.5% 2.03e-1 8.23
Llama3.2 3B +DE-MARK 8.0% 1.5% 1.0% 5.00e-1 8.28
TPR@FPR=
Concept Probing (short answer) 10% ↓ 1% ↓ 0.1% ↓ Median p-value ↑ F1 ↑
Llama3.1 8B +Watermark 19.0% 11.5% 5.0% 2.82e-1 42.55
Llama3.1 8B +DE-MARK 9.5% 3.0% 0.0% 5.00e-1 45.31
Mistral 7B +Watermark 41.5% 11.5% 2.0% 1.59e-1 35.95
Mistral 7B +DE-MARK 21.5% 0.5% 0.0% 5.00e-1 37.8
Llama3.2 3B +Watermark 14.0% 4.5% 1.5% 2.82e-1 33.50
Llama3.2 3B +DE-MARK 5.0% 2.5% 0.0% 5.00e-1 39.13
TPR@FPR=
Long-form QA (long answer) 10% ↓ 1% ↓ 0.1% ↓ Median p-value ↑ Rouge-L ↑
Llama3.1 8B +Watermark 98.5% 96.5% 91.0% 1.07e-10 23.99
Llama3.1 8B +DE-MARK 3.0% 0.5% 0.5% 1.14e-2 23.75
Mistral 7B +Watermark 99.0% 95.5% 88.5% 7.66e-9 23.20
Mistral 7B +DE-MARK 13.0% 3.0% 1.5% 3.23e-2 23.18
Llama3.2 3B +Watermark 99.5% 98.5% 93.0% 2.25e-10 23.74
Llama3.2 3B +DE-MARK 8.5% 2.5% 0.0% 5.30e-2 23.94
@@
Appendix C Ablation study
@@
[…]
0.1$\downarrow$ 0.01$\downarrow$ 0.001% $\downarrow$
@@
Appendix D Additional experimental results
@@
Additionally, we evaluated the effectiveness of our method against two n-gram-based distortion-free watermarking schemes, $\gamma$-reweight (Hu et al., 2023) and DiPmark (Wu et al., 2023b), as shown in Table 11 and Table 12. The results demonstrate that our proposed De-mark effectively removes both types of watermarks.
@@
TPR@FPR= median $\downarrow$ GPT $\uparrow$
0.1$\uparrow$ 0.01$\uparrow$ 0.001$\uparrow$
@@
Table 11: Watermark removal results against $\gamma$-reweight
MMW Book Report Dolly CW
TPR@FPR= median p-value $\uparrow$ GPT score $\uparrow$ TPR@FPR= median p-value $\uparrow$ GPT score $\uparrow$
0.1$\downarrow$ 0.01$\downarrow$ 0.001$\downarrow$ 0.1$\downarrow$ 0.01$\downarrow$ 0.001$\downarrow$
Llama3.1 8B +Watermark 44% 15% 7% 2.48e-3 94.72 42% 36% 29% 3.08e-3 93.70
+De-mark 5% 1% 1% 9.01e-2 94.51 25% 19% 17% 6.95e-2 92.45
Llama3.2 3B +Watermark 42% 20% 11% 2.48e-3 94.51 34% 23% 16% 7.75e-3 92.25
+De-mark 15% 7% 4% 2.94e-2 94.11 25% 17% 13% 5.29e-2 91.58
Table 12: Watermark removal results against DiPmark
MMW Book Report Dolly CW
TPR@FPR= median p-value $\uparrow$ GPT score $\uparrow$ TPR@FPR= median p-value $\uparrow$ GPT score $\uparrow$
1$\downarrow$ 0.1$\downarrow$ 0.01$\downarrow$ 1$\downarrow$ 0.1$\downarrow$ 0.01$\downarrow$
Llama3.1 8B +Watermark 44% 22% 11% 1.55e-2 94.58 43% 25% 17% 1.55e-2 92.68
+De-mark 3% 2% 0% 6.53e-1 94.75 5% 2% 0% 6.18e-1 92.34
Llama3.2 3B +Watermark 43% 27% 13% 1.85e-2 94.24 39% 23% 18% 2.15e-2 91.98
+De-mark 2% 1% 1% 4.46e-1 93.78 4% 3% 1% 5.48e-1 91.79
Appendix E Proof of Theorem 4.1
@@
Assume w.l.o.g. $T_{i}$ in the green list and $T_{j}$ in the red list. Based on Equation 1, the watermarked distribution $P_{W}(T_{i}|\bm{x},T_{i},T_{j})$ and $P_{W}(T_{j}|\bm{x},T_{i},T_{j})$ should be $P_{W}(T_{i}|\bm{x},T_{i},T_{j})=\frac{(0.5+\epsilon)e^{\delta}}{(0.5+\epsilon)e^{\delta}+(0.5-\epsilon)}$ and $P_{W}(T_{j}|\bm{x},T_{i},T_{j})=\frac{(0.5-\epsilon)}{(0.5+\epsilon)e^{\delta}+(0.5-\epsilon)}$, where $\epsilon$ is the noise. In this case,
@@
Appendix F Proof of Theorem 4.2
@@
Appendix G Broader Impact
Machine learning models are transforming various sectors by improving efficiencies and tackling complex problems. Despite these beneficial impacts, there are valid concerns regarding the integrity and security of machine learning implementations (Wang et al., 2023d, c; Wu et al., 2022, 2023a, 2023c; Mai et al., 2023b, a; Guo et al., 2024). In this context, watermarking has become a critical tool. It not only verifies the authenticity and ownership of digital media but also aids in identifying AI-generated content.